During the past few years, Harry Davis Industries (HDI) has been constrained by high cost of capital to make many capital investments. Recently, though, capital costs have been declining and the company has decided to look seriously at a major expansion program that had been proposed by the marketing department. Assume that you are an assistant to the CFO. Your first task is estimate HDI’s cost of capital. The CFO has provided you with the following data, which is considered to your task:

  1. The current price of HDI’s 12% coupon , seminal annual, non-callable bonds with 15 years to maturity is $1153.72. HDI does not use short-term interest bearing debt on a permanent basis. New bonds would be privately placed with no flotation costs.
  2. The current price of HDI’s 10%, $100 par value, quarterly dividend, perpetual preferred stock is $113.10. HDI would incur flotation cost of $2.00 per share.
  3. HDI’s common stock is currently selling for $50 per share. Its last dividend (d0) was $4.19, and dividends are expected to grow at a constant rate of 5% in the foreseeable future. HDI’s beta is 1.2; the yield on T-Bonds is 7%; and the market risk premium is estimated to be 6%. For the bond-yield-plus-risk-premium approach, the firm uses a 4% point risk premium.
  4. HDI’s target capital structure:30% long-term, 10% pref. stock, and 60% common equity.
  5. The firm’s tax rate is 40%

To structure the task somewhat, CFO has asked you to answer the following questions:

1.   The after-tax cost of HDI’s debt, Kd is ___________

2.   The firm’s cost of preferred stock, Kp, is __________

3.   Using CAPM the firm’s cost of equity, Ke is _______

4.   The estimated cost of equity using Discounted Cash Flow (DCF) model is ______

5.   The cost of equity, Ke, based on the bond-yield-plus-risk-premium method is ____

6.   The overall cost of equity, Ke is __________

7.   The weighted average cost of capital (WACC) is _________

8.   HDI estimates that if it issues new common stock, the flotation cost will be 15%.   HDI incorporates the flotation costs into the DCF         approach. The estimated cost of newly issued common stock, taking into account the flotation cost is _________

9. Suppose HDI has historically earned 15% on equity (ROE) and retained 35% of earnings, and investors expect this situation to continue in the future. How could you use this information to estimate the future dividend growth rate, and what growth rate would        you get? Is this consistent with the 5% growth rate given?

HDI is interested in establishing a new division, which will focus primarily on developing new internet-based projects. In trying to determine the cost of capital for this new division, you discover that stand-alone firms involved in similar projects have on average the following characteristics:

- Capital structure of 10% debt and 90% equity.

- Cost of debt of 12% and beta of 1.7.

10. Given this information, your estimate of the division’s cost capital is _______
Answer #1
1.The after-tax cost of HDI’s debt, Kd is:
Using the formula to find Present Value of a bond,
PV of a bond=PV of its future coupons+PV of Fv to be recd. At maturity
ie.PV of a bond=(Pmt.*(1-(1+r)^-n)/r)+(FV/(1+r)^n)
where,
PV= Price/Present market value of the bond.--here $ 1153.72
Pmt.=periodic (here,semi-annual) coupon on the bond=1000*12%/2= 60
r=   ? the yield on the bond-- to find out
n=no.of periodic payments=15*2= 30 semi-annual periods
FV= Face value, $ 1000
Applying these values, in the above formula,
1153.72=(60*(1-(1+r)^-30)/r)+(1000/(1+r)^30)
Solving for r, we get the before-tax semi-annual yield ,
r= 5%
Before-tax Annual r=(1+0.05)^2-1=
10.25%
So, after tax annual cost of debt, Kd=
10.25%*(1-40%)=
6.15%
2.The firm’s cost of preferred stock, Kp, is
Annual dividend/Net Proceeds of issue
ie.Annual dividend/(current market price-Flotation cost)per share
(10%*100)/(113.1-2)=
9%
3.Using CAPM the firm’s cost of equity, Ke is
Ke=RFR+(Beta*Market-risk premium)
ie. Ke=7%+(1.2*6%)=
14.2%
4. The estimated cost of equity using Discounted Cash Flow (DCF) model is
Using DCF model formula for cost of equity,
Ke=(D1/P0)+g
where, Ke=   ? cost of equity-- to be found out
D1=Next dividend,ie.D0*(1+g)=4.19*(1+5%)= 4.3995
P0= Current market price = $ 50
So, applying the values in the formula,
Ke=(4.3995/50)+5%=
13.80%
5.Cost of equity, Ke, based on the bond-yield-plus-risk-premium method is
6.15%+4%=
10.15%
6. Overall cost of Equity
Taking the average cost for equity (14.2%+13.8%+10.15%)/3= 12.72%
7.Weighted average cost of capital (WACC) is
WACC=(Wd*Kd)+(Wp*Kp)+(We*Ke)
ie. WACC=(30%*6.15%)+(10%*9%)+(60%*12.72%)=
10.38%
8.The estimated cost of newly issued common stock, taking into account the flotation cost is
Ke=(4.3995/(50*(1-15%)))+5%=
15.35%
9...Continued growth rate or
Sustained growth rate= ROE*RR
ie. 15%*35%=
5.25%
As RR depends on dividend pay-out,
ie. Increased dividend payout decreases the RR
As the investors expect a SGR of 5.25%, care should be taken to design a dividend pay-out ratio , that enables to retain that %age which when applied to ROE works out to 5.25%.
That way, future dividends can be estimated within this framework.
NO.
5.25%> 5%
10.Given this information, estimate of the division’s cost capital is :
Cost of equity with beta of 1.7
ie. Ke=7%+(1.7*6%)=
17.20%
WACC(new division)= (Wd*Kd.(1-Tax rate))+(We*Ke)
(10%*12%*(1-40%))+(90%*17.2%)=
16.2%